OpenStudy (cicilybailey14):

Help please :D Graph the slope and y-intercept Slope: 1/2 Y-intercept: -4

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

y = 1/2x + -4

OpenStudy (anonymous):

Great graphing calculator.

OpenStudy (anonymous):

That is the result after plugging those into y = mx + b where m is the slope and by is the y-intercept.

OpenStudy (anonymous):

@Cicilybailey14 Plug y = 1/2x + -4 into https://www.desmos.com/calculator and see what you get.

OpenStudy (igreen):

y = mx + b Where m is the slope, and b is the y-intercept. So here we can plug in 1/2 for 'm' and -4 for 'b', giving us: y = 1/2x - 4

OpenStudy (anonymous):

y = mx + b Slope = m y-intercept = b y = \(\dfrac{1}{2}\)x - 4.

OpenStudy (cicilybailey14):

How do you work the site lol